OpenStudy (anonymous):

what is -12 divided by 2/3?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Use a calculator.. lol

OpenStudy (anonymous):

how do i do fractions on a calculator??

OpenStudy (trojanpoem):

Dude, -12 / 2/3 = -12 * 3/2 = 6*3 = - 18

OpenStudy (anonymous):

-18

OpenStudy (trojanpoem):

how do i do fractions on a calculator?? << Use calculator and I will guarantee that you won't do any math problem for the rest of your life.